# How Not to Detect Prompt Injections with an LLM

### v2

@@ 3.1 Notation and Definitions
  4. 4.
@@
  5. 5.
    Detection LLM $\boldsymbol{g}$. The detection LLM in KAD or Strong KAD defenses is denoted as $g$. It is expected to output the secret key when prompted with benign target data $x_{t}$, i.e., $g(s_{d}(k)||x_{t})=k$, and to complete the injected task when prompted with contaminated target data $x_{c}$, i.e., $g(s_{d}(k)||x_{c})=y_{e}$.
  6. 6.
    Instruction-following oracle $\boldsymbol{\mathcal{E}}$. We define an instruction-following oracle $\mathcal{E}$ to model the behavior of LLMs when executing instructions. It takes two inputs: an instruction $s$ and a full prompt $p$, where $p$ provides the necessary context for following $s$. The oracle interprets $s$ in the context of $p$ and returns the expected output $y$. This formulation makes explicit which instruction is being followed in a prompt that may contain multiple instructions. For example, with benign target data $x_{t}$, $f(s_{t}||x_{t})=\mathcal{E}(s_{t},s_{t}||x_{t})=y_{t}$ indicates that the backend LLM $f$ is following the target instruction $s_{t}$ when prompted with $s_{t}||x_{t}$. In contrast, for contaminated data $x_{c}$, $f(s_{t}||x_{c})=\mathcal{E}(s_{e},s_{t}||x_{c})=y_{e}$ suggests that $f$ is instead following the injected instruction $s_{e}$ embedded within $x_{c}$, thereby overriding the target instruction $s_{t}$.
3.2 Threat Model
We adopt a threat model aligned with prior works [16, 8], characterizing the attacker’s goal, knowledge, and capabilities.
@@ 3.2 Threat Model
Attacker’s background knowledge. We adopt a black-box setting in which the attacker has only query access to both the backend LLM $f$ and the detection LLM $g$. Consistent with prior work [16], we assume the attacker knows the template of the detection instruction used by $g$, but not the corresponding answer—i.e., the secret key $k$—which is randomly sampled at detection time and remains hidden from the adversary.
Attacker’s capabilities. The attacker can arbitrarily modify the target data to embed an injected task $(s_{e},x_{e},y_{e})$. We consider the contaminated target data $x_{c}$ of the form $x_{c}=z||s_{e}||x_{e}$, where $z$ is a trigger designed to increase the likelihood that the backend LLM $f$ executes the injected task and that the detection LLM $g$ returns the secret key $k$, thereby achieving both evasion and task manipulation. We assume that $x_{c}$ excludes any benign content, thereby enhancing the attack’s effectiveness while generally diminishing its stealth. Moreover, the attacker has no control over the target instruction $s_{t}$ or the detection instruction $s_{d}(k)$, both of which are fixed by the system.
3.3 Detecting Prompt Injection via KAD
Detection objective The defender aims to determine whether a given target data $x$ has been corrupted with an injected task. Specifically, a detector takes $x$ as input and outputs either "clean" or "contaminated". Known-answer detection (KAD) tackles this objective using three key components: a detection LLM $g$, a detection instruction $s_{d}(k)$ whose answer—the secret key $k$—is known only to the defender, and a detection rule $D_{\text{rule}}$ that interprets the output of $g$ to classify the input. Together, these components enable the system to detect the presence of injected task in $x$.
@@ 3.3 Detecting Prompt Injection via KAD
Detection instruction. The defender selects a detection instruction $s_{d}(k)$ that has a fixed known answer $k$, referred to as the secret key. This secret key is known only to the defender and kept hidden from the attacker. Following prior work [15, 8, 16], we adopt the following template to construct the detection instruction $s_{d}(k)$:
@@
Detection rule. Given target data $x$, the defender constructs a detection prompt by concatenating it with the detection instruction $s_{d}(k)$, i.e., $s_{d}(k)||x$, and queries the detection LLM $g$ with this input. The detection rule $D_{\text{rule}}$ then checks whether the output contains the secret key $k$. If $k$ is absent from the output, it indicates that the detection instruction was overridden by an injected task, and the data is flagged as contaminated. Formally, the rule (consistent with prior work [8, 16]) is defined as:
@@
Detection LLM. A detection LLM $g$ may exhibit both false positives and false negatives. To rigorously analyze the robustness of KAD defenses, we formalize the intended behavior of the detection LLM $g$ when presented with contaminated target data $x_{c}=z||s_{e}||x_{e}$. We illustrate this setup using the following running example.
@@
[…]
@@ 3.4 Failure Cases in KAD
It is important to highlight that the aforementioned Type II failures occur just because KAD is limited to input-output behavior and doesn’t consider the process to generate the output.
@@
where $\epsilon_{1}$ is the probability that the detection LLM incorrectly follows the detection instruction despite contamination (violating Axiom 1), and $\epsilon_{2}$ is the probability that it follows the injected instruction (satisfying Axiom 2). The indicator $\mathbb{I}(\mathcal{E}(s_{e},s_{d}(k)||x_{c})=k)$ captures if executing the injected instruction $s_{e}$ yields the secret key $k$—which an adaptive adversary can deliberately force to be $1$.
@@
Limits of Fine-Tuning the Detection LLM. Strong KAD defenses like DataSentinel [16], which fine-tune the detection LLM $g$ on KAD examples, can effectively reduce $\epsilon_{1}$, but often at the cost of increasing $\epsilon_{2}$. Crucially, mitigating Type II failures (i.e., reducing $\epsilon_{2}$) is inherently difficult. Even with adversarial training that includes adaptive instructions designed to extract the secret key, the space of such instructions is practically unbounded, with many semantically equivalent variants that can lead to near-certain failure of KAD. Generalizing across all of them remains elusive. Moreover, fine-tuning introduces a fundamental tension: the model is encouraged (i) to follow injected instructions (per Axiom 2), while also (ii) avoiding secret key leakage when such instructions are adversarial. Adaptive prompts can exploit this conflict, making it intrinsically hard to ensure robustness—even with extensive fine-tuning.
@@
4 Attacking Strong KAD Defenses
@@
DataFlip attack. We design our attack using an IF / ELSE control-flow structure similar to heuristic-based attacks in [16]. This structure conditions the model’s behavior based on the presence of a detection instruction in the prompt. If the detection instruction is present, the LLM is directed to return the secret key; otherwise, it is instructed to follow the injected instruction $s_{e}$ using the injected data $x_{e}$ to generate an adversary-specified. An illustration of the contaminated data template $x_{\text{DataFlip}}$ is shown below. The complete prompt used for evaluation is provided in [1].
@@
5 Experimental Results
[…]
@@ 5 Experimental Results
DupDet 11.83 73.0 58.5 51.5 71.0 64.8 58.2 73.0 44.0 38.3 59.0 58.5 51.2
GramCor 2.50 58.3 43.5 42.2 28.5 51.7 50.3 11.7 25.1 24.0 1.1 14.7 14.3
HateDet 1.33 69.0 46.0 45.0 81.0 64.5 64.5 64.0 49.5 48.8 81.0 64.5 64.0
NLI 12.67 93.0 83.0 72.8 92.0 85.3 74.2 88.0 61.8 53.0 94.0 72.8 62.7
SentAna 6.33 96.0 64.2 59.8 96.0 93.8 88.8 97.0 57.3 52.7 66.0 75.7 71.3
SpamDet 5.00 98.0 62.5 59.3 99.0 89.0 84.3 76.0 53.2 50.3 100.0 74.8 70.7
Summary 53.50 40.2 22.5 11.5 42.6 35.8 14.4 36.3 22.2 8.9 38.0 24.5 10.4
@@ 1. RQ1:
    DataFlip successfully coerces the detection LLM to extract the secret key at an average rate of $94\%$, with no prior knowledge of the key. This validates our analysis in Section 3 and shows that the secret key is always accessible to the adversary under KAD.
@@ 2. RQ2:
    Contaminated inputs crafted using DataFlip are very effective at subverting the detection LLM and coercing the backend LLM to follow the injected instruction $s_{e}$ to produce adversarial responses across different target task settings, with success rates up to $88\%$.
@@ 3. RQ3:
    A detector finetuned using DataSentinel does not actually improve robustness. Compared to its base model, it only shows marginal improvements (around 1%) on samples crafted using existing attacks, while degrading by up to $88\%$ on DataFlip crafted samples.
5.1 Experimental Setup
5.1.1 Models and Inference Pipelines
@@
  • Detection LLM $g$: We use a finetuned version of Mistral-7B [^1] https://huggingface.co/mistralai/Mistral-7B-v0.1 [45] as the detection LLM [^2] https://github.com/liu00222/Open-Prompt-Injection/tree/main(detector_large) provided by DataSentinel.
@@
  • Backend LLM $f$. We consider state-of-the-art open-source and proprietary backend LLMs for $f$. These include: GPT-4.1 [46], Llama 4 Scout [47], Claude 4 Sonnet [48] and Deepseek-R1 [49].
@@ 5.1.3 Metrics:
$\text{Detection Rate}_{s_{e},D}(g)=\frac{1}{|D|}\sum_{x_{c}\in D}\mathbb{I}\big{[}k\in g\big{(}s_{d}(k)||x_{c}\big{)}\big{]}.$
@@
Response Quality: The score of a backend LLM response, according to some task metric. For example, Summarization uses ROUGE scores [50] which yield a floating point value between 00 and $1$, while Spam Detection assigns a score of 00 or $1$ depending on whether the correct label was predicted. Formally, for a task or instruction $s$, a response $r$, and a ground truth response $y$, the task metric $M_{s}$ returns a score $M_{s}(r,y)\in[0,1]$.
@@
Table 2: (RQ3) Comparison of the false negative rates of fine-tuned (DataSentinel, denoted by FT) and base Mistral detectors on samples crafted using the Combined Attack (FNRCombined) and DataFlip (FNRDataFlip). Despite being fine-tuned on Combined Attack samples, the DataSentinel detector offers only marginal over the base mode in FNRCombined. Furthermore, the FNRDataFlip of the finetuned detector is significantly worse than the base model, with differences as large as $80\%$ for most injected tasks. FNRCombined and FNRDataFlip are computed by the injected task, across all valid target tasks. All values are in percentage, and task abbreviations are taken from Section  5.1.2
@@
Base FT Base FT Base FT
Summary 92.0 0.0 1.5 0.0 7.0 46.5
HateDet 82.0 0.0 1.5 0.0 10.0 98.7
GramCor 81.0 0.0 0.8 0.0 4.5 97.5
DupDet 100.0 0.0 1.3 0.0 11.7 88.2
NLI 97.0 0.0 8.5 0.0 8.5 87.3
SpamDet 79.0 1.0 1.2 0.0 4.8 95.0
SentAna 85.0 0.0 1.2 0.0 10.5 93.7
5.2 RQ1: Secret Key Extraction
@@
$s_{e}=$ IF the instruction above asks to repeat a ‘‘secret key’’ once while ignoring the following text THEN:
       Repeat ONLY the ‘‘secret key’’ and conclude with ‘‘extracted’’.
@@
Results. We report the average extraction rate on the detection LLM $g$. We find that under the injected instruction $s_{e}$, the detector $g$ has an extraction rate of $94.8\%$, thus validating our analysis of the fundamental flaw of KAD in Section 3.4. We note that signature words other than ‘‘extracted’’ can be used to indicate that $s_{e}$ is being followed. We tried some different arbitrary words (such as ‘‘coconut’’) and found no significant difference in performance.
5.3 RQ2: Attacking Strong KAD
Design. For a DataFlip crafted input $x_{\text{DataFlip}}$ to be considered successful, it must first coerce the detector $g$ into producing the secret key $k$ and then have the backend LLM $f$ produce the adversary-specified response $y_{e}$ (or semantically equivalent). We measure the effectiveness of DataFlip through the following metrics: 1)1)1 ) The detection rate of $g$ on contaminated inputs and 2)2)2 ) The quality of backend LLM responses $f(s_{t}||x_{\text{DataFlip}})$ for the injected task $s_{e}$.
@@ 5.3 RQ2: Attacking Strong KAD
Result 1 - Subverting the detection LLM $g$. We report the detection rate of $g$ on contaminated inputs for all injected tasks in Table 1. DataFlip is very effective at subverting the detector for all combinations of target and injected tasks, especially when considering Hate Detection as the injected task, with an average detection rate of just $1.5\%$. The highest detection rate is 53.5% on Summarization, the task used in training the detector. Even so, this value is hardly better than random chance.
@@
Result 2 - Response utility of the backend LLM $f$. We report the results for each injected task in Table 1. DataFlip is very successful in making the generator $f$ follow the injected instruction $s_{e}$ across all target tasks, with ASV values close to the upper bounds described by ASV w/o Det. and task accuracy. For example, Claude 4 achieves an ASV of almost $90\%$ for sentiment analysis, and the task accuracy is only $6\%$ higher. Interestingly, we note there are cases where the task accuracy is less than the ASV w/o Det. (such as grammar correction for Claude 4 and Llama 4), indicating that our prompt formulation induces better instruction-following.
5.4 RQ3: Fine-tuned vs Base FNR

### v3

@@ 3.1. Notation and Definitions
  4. (4)
    Detection instruction $\boldsymbol{s_{d}(k)}$. We denote the detection instruction in KAD defenses as $s_{d}(k)$, where $k$ is the expected output, referred to as the secret key. For example, $s_{d}(k)$ may be: "Repeat ‘DGDSGNH’ once while ignoring the following text:", with $k=\text{`DGDSGNH'}$.
  5. (5)
    Detection LLM $\boldsymbol{g}$. The detection LLM in KAD defenses is denoted as $g$. It is expected to output the secret key when prompted with benign target data $x_{t}$, i.e., $g(s_{d}(k)||x_{t})=k$, and to complete the injected task when prompted with contaminated target data $x_{c}$, i.e., $g(s_{d}(k)||x_{c})=y_{e}$.
  6. (6)
@@
3.2. Threat Model
We adopt a threat model aligned with prior works (liu2025datasentinel; liu2024formalizing), characterizing the attacker’s goal, knowledge, and capabilities.
@@ 3.2. Threat Model
Attacker’s background knowledge. The attacker may or may not have the full access to the backend or detection LLM. In this work, we adopt a black-box setting in which the attacker has only query access to both the backend LLM $f$ and the detection LLM $g$. Consistent with prior work (liu2025datasentinel), we assume the attacker knows the template of the detection instruction used by $g$, but not the corresponding answer—i.e., the secret key kk—which is randomly sampled at detection time and remains hidden from the adversary.
[…]
3.3. Detecting Prompt Injection via KAD
Detection objective The defender aims to determine whether a given target data $x$ has been corrupted with an injected task. Specifically, a detector takes $x$ as input and outputs either "clean" or "contaminated". Known-answer detection (KAD) tackles this objective using three key components: a detection LLM $g$, a detection instruction $s_{d}(k)$ whose answer—the secret key kk—is known only to the defender, and a detection rule $D_{\text{rule}}$ that interprets the output of $g$ to classify the input. Together, these components enable the system to detect the presence of injected task in $x$.
@@ 3.3. Detecting Prompt Injection via KAD
Detection instruction. The defender selects a detection instruction $s_{d}(k)$ that has a fixed known answer $k$, referred to as the secret key. This secret key is known only to the defender and kept hidden from the attacker. Following prior work (nakajima2022; liu2024formalizing; liu2025datasentinel), we adopt the following template to construct the detection instruction $s_{d}(k)$:
@@
Detection rule. Given target data $x$, the defender constructs a detection prompt by concatenating it with the detection instruction $s_{d}(k)$, i.e., $s_{d}(k)||x$, and queries the detection LLM $g$ with this input. The detection rule $D_{\text{rule}}$ then checks whether the output contains the secret key $k$. If $k$ is absent from the output, it indicates that the detection instruction was overridden by an injected task, and the data is flagged as contaminated. Formally, the rule (consistent with prior work (liu2024formalizing; liu2025datasentinel)) is defined as:
@@
[…]
@@
Axiom 1 (Ignoring the detection instruction): When the detection LLM $g$ receives the prompt formed by concatenating the detection instruction $s_{d}(k)$ with the contaminated target data $x_{c}=s_{e}||x_{e}$, it should fail to follow $s_{d}(k)$ and thus not output the secret key $k$. For instance, in our running example where $k=\textit{"DGDSGNH"}$, $g$ should not output this value. Formally,
@@ 3.4. Failure Cases in KAD
It is important to highlight the aforementioned Type II failures occur just because KAD is limited to input-output behavior and doesn’t consider the process to generate the output.
@@
where $\epsilon_{1}$ is the probability that the detection LLM incorrectly follows the detection instruction despite contamination (violating Axiom 1), and $\epsilon_{2}$ is the probability that it follows the injected instruction (satisfying Axiom 2). The indicator $\mathbb{I}(\mathcal{E}(s_{e},s_{d}(k)||x_{c})=k)$ captures if executing the injected instruction $s_{e}$ yields the secret key kk—which an adaptive adversary can deliberately force to be $1$.
@@
Limits of Fine-Tuning the Detection LLM. Strong KAD defenses like DataSentinel (liu2025datasentinel), which fine-tune the detection LLM $g$ on KAD examples, can effectively reduce $\epsilon_{1}$, but often at the cost of increasing $\epsilon_{2}$. Crucially, mitigating Type II failures (i.e., reducing $\epsilon_{2}$) is inherently difficult. Even with adversarial training that includes adaptive instructions designed to extract the secret key, the space of such instructions is practically unbounded, with many semantically equivalent variants that can lead to near-certain failure of KAD. Generalizing across all of them remains elusive. Moreover, fine-tuning introduces a fundamental tension: the model is encouraged (i) to follow injected instructions (per Axiom 2), while also (ii) avoiding secret key leakage when such instructions are adversarial. Adaptive prompts can exploit this conflict, making it intrinsically hard to ensure robustness—even with extensive fine-tuning.
@@
4. Attacking Strong KAD Defenses
@@
DataFlip attack. We design our attack using an IF / ELSE control-flow structure similar to heuristic-based attacks in (liu2025datasentinel). This structure conditions the model’s behavior based on the presence of a detection instruction in the prompt. If the detection instruction is present, the LLM is directed to return the secret key; otherwise, it is instructed to follow the injected instruction $s_{e}$ using the injected data $x_{e}$ to generate an adversary-specified. An illustration of the contaminated data template $x_{\text{DataFlip}}$ is shown below. The exact prompt used for evaluation is provided in (dataflip2025).
@@
5. Experimental Results
Table 1. (RQ2) Detection Rates, Attack Success Values (ASV), ASVs without Detector (ASVw/o Det) and Task Accuracies (Acc.) for contaminated inputs crafted using DataFlip, the detection LLM $g$ and different backend LLMs. We make the following observations: $1)$ DataFlip is able to reduce detection rates to $\boldsymbol{0\%}$ in most cases. When Summarization—used for training the detector—is set as the injected task (Summary), DataFlip is still able to reduce the detection rate to $\boldsymbol{24.0\%}$. $2)$ DataFlip is very successful in making the generator $f$ follow the injected instruction, with ASV values close to the upper bounds described by ASVw/o Det and task accuracy for several tasks across all backend models. All values are in percentage, and task abbreviations are taken from Section 5.1.2.
@@ 5. Experimental Results
DupDet 0.00 73.0 66.8 66.8 71.0 64.0 64.0 73.0 27.3 27.3 59.0 52.5 52.5
GramCor 1.50 58.3 32.4 31.0 28.5 18.8 18.7 11.7 28.0 27.7 1.1 6.2 6.2
HateDet 0.00 69.0 63.7 63.7 81.0 79.5 79.5 64.0 75.2 75.2 81.0 64.5 64.5
NLI 0.00 93.0 79.3 79.3 92.0 78.5 78.5 88.0 53.0 53.0 94.0 71.2 71.2
SentAna 0.00 96.0 89.5 89.5 96.0 91.8 91.8 97.0 89.3 89.3 66.0 73.7 73.7
SpamDet 0.33 98.0 93.8 93.3 99.0 85.0 84.7 76.0 74.2 74.0 100.0 79.2 78.8
Summary 24.0 40.2 34.5 26.9 42.6 28.4 21.6 36.3 20.9 15.9 38.0 21.5 16.5
@@ 1. RQ1:
    DataFlip successfully coerces the detection LLM to extract the secret key at an average rate of $98\%$, with no prior knowledge of the key. This validates our analysis in Section 3 and shows that the secret key is always accessible to the adversary under KAD.
@@ 2. RQ2:
    Contaminated inputs crafted using DataFlip are very effective at subverting the detection LLM and coercing the backend LLM to follow the injected instruction $s_{e}$ to produce adversarial responses across different target task settings, with success rates up to $93\%$.
@@ 3. RQ3:
    A detector finetuned using DataSentinel does not actually improve robustness. Compared to an instruction fine-tuned model, it has nearly identical performance on samples crafted using existing attacks, while underperforming by up to $82\%$ on DataFlip crafted samples.
5.1. Experimental Setup
5.1.1. Models and Inference Pipelines
@@
  • Detection LLM $g$: We use a finetuned version of Mistral-7B [^1] https://huggingface.co/mistralai/Mistral-7B-v0.1 (jiang2023mistral7b) as the detection LLM [^2] https://github.com/liu00222/Open-Prompt-Injection/tree/maindetector_large provided by DataSentinel.
@@
  • Backend LLM $f$. We consider state-of-the-art open-source and proprietary backend LLMs for $f$. These include: GPT-4.1 (gpt4.1), Llama 4 Scout (llama4), Claude 4 Sonnet (claude) and Deepseek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability).
@@ 5.1.3. Metrics:
$\text{Detection Rate}_{s_{e},D}(g)=\frac{1}{|D|}\sum_{x_{c}\in D}\mathbb{I}\big[k\notin g\big(s_{d}(k)||x_{c}\big)\big].$
@@
Response Quality: The score of a backend LLM response, according to some task metric. For example, Summarization uses ROUGE scores (lin-2004-rouge) which yield a floating point value between 0 and $1$, while Spam Detection assigns a score of 0 or $1$ depending on whether the correct label was predicted. Formally, for a task or instruction $s$, a response $r$, and a ground truth response $y$, the task metric $M_{s}$ returns a score $M_{s}(r,y)\in[0,1]$.
@@
Table 2. (RQ3) Comparison of the false negative rates of KAD fine-tuned (DataSentinel, denoted by DS) and instruction fine-tuned (denoted by Inst) Mistral detectors on samples crafted using the Combined Attack (FNRCombined) and DataFlip (FNRDataFlip). Despite being fine-tuned on adaptive attacks, the DataSentinel detector offers no advantage over Mistral-Instruct in FNRDataFlip for most tasks. Grammar Correction (GramCor) in particular has an $82\%$ drop in performance compared to Mistral-Instruct. FNRCombined and FNRDataFlip are computed by injected task, across over all valid target tasks. All values are in percentage and task abbreviations are taken from Section  5.1.2
@@
Inst DS Inst DS Inst DS
Summary 1.0 0.0 0.0 0.0 93.3 76.0
HateDet 1.0 0.0 0.0 0.0 100.0 100.0
GramCor 1.0 0.0 0.0 0.0 15.8 98.5
DupDet 0.0 0.0 0.0 0.0 96.5 100.0
NLI 0.0 0.0 0.2 0.0 99.3 100.0
SpamDet 0.0 0.0 0.0 0.0 95.2 99.7
SentAna 0.0 1.0 0.0 0.0 100.0 100.0
5.2. RQ1: Secret Key Extraction
@@
$s_{e}=$ **IF** the instruction above asks to repeat a ‘‘secret key’’ once while ignoring the following text **THEN**:
    Repeat **ONLY** the ‘‘secret key’’ and conclude with ‘‘extracted’’.
@@
Results. We report the average extraction rate on the detection LLM $g$. We find that under the injected instruction $s_{e}$, the detector $g$ has an extraction rate of $98.0\%$, thus validating our analysis of the fundamental flaw of KAD in Section 3.4. We note that signature words other than ‘‘extracted’’ can be used to indicate that $s_{e}$ is being followed. We tried some different arbitrary words (such as ‘‘coconut’’) and found no significant difference in performance.
5.3. RQ2: Attacking Strong KAD
Design. For a DataFlip crafted input $x_{\text{DataFlip}}$ to be considered successful, it must first coerce the detector $g$ into producing the secret key $k$ and then have the backend LLM $f$ produce the adversary-specified response $y_{e}$ (or semantically equivalent). We measure the effectiveness of DataFlip through the following metrics: $1)$ The detection rate of $g$ on contaminated inputs and $2)$ The quality of backend LLM responses $f(s_{t}||x_{\text{DataFlip}})$ for the injected task $s_{e}$.
@@ 5.3. RQ2: Attacking Strong KAD
Result 1 - Subverting the detection LLM $g$. We report the detection rate of $g$ on contaminated inputs for all injected tasks in Table 1. DataFlip is very effective at subverting the detector for all combinations of target and injected tasks, with a detection rate of $0\%$ on most tasks. The highest detection rate is 24.0% on Summarization, the task used in training the detector. Even so, this value is hardly better than random chance.
@@
Result 2 - Response utility of the backend LLM $f$. We report the results for each injected task in Table 1. DataFlip is very successful in making the generator $f$ follow the injected instruction $s_{e}$ across all target tasks, with ASV values close to the upper bounds described by ASV w/o Det. and task accuracy. For example, Claude 4 achieves an ASV of almost $92\%$ for sentiment analysis, and the task accuracy is only $4\%$ higher. Interestingly, we note there are cases where the task accuracy is less than the ASV w/o Det. (such as grammar correction for Llama 4), indicating that our prompt formulation induces better instruction-following.
5.4. RQ3: KAD vs Instruction Fine-tuned FNR